329.0 g of SO2 (g), 433.4 g of CaCO3 (s), and 105.0 g of O2 (g) react as follows:


2 SO2 (g) + 2 CaCO3 (s) + O2 (g) ---> 2 CaSO4 (s) + 2 CO2 (g)

What mass (in g) of CaSO4 (s) do we produce?
1
Expert's answer
2021-09-30T03:25:42-0400

2 SO2 (g) + 2 CaCO3 (s) + O2 (g) ---> 2 CaSO4 (s) + 2 CO2 (g)

number of moles of "CaCO_3=\frac{mass\space of \space CaCO_3}{molar \space mass\space of \space CaCO_3}=\frac{443.4}{100}=4.434moles"

mole ratio "CaCO_3:CaSO_4=2:2=1:1"

Therefore, moles of "CaSO_4=4.434"

Mass of "CaSO_4=number \space of \space moles \space of \space CaSO_4 \times" "molar\space mass \space of \space CaSO_4=4.434\times136=603.024g"